Make New Entry,
Make Followup Entry
User name Sirish
Log entry time 21:19:36 on November17,2008
Entry number 249436
Established beam through Compton Chicane. No beam loss trips.
Photon Background (blue trace) rate is 35 kHz/uA with normal settings, while our goal is ~200 Hz/uA. Turning off Hall B drops the rates to 2 kHz/uA. Closing injector slits to 15.2 drops the rate further to 400 Hz/uA. This means the injector settings are not optimal and need tweaking to clean up Hall A beam.
We will keep the beam in the Chicane with cavity off to debug the e-detector readout. The e-detector must stay in garage location until we have low background tune.